# Addressing Sample Inefficiency in Multi-View Representation Learning

### v1

@@ 4.2 Corollary 1: Low-dimensional projectors are sufficient
While BarlowTwins and VICReg frameworks have advocated using high-dimensional projectors to facilitate good feature learning on Imagenet, our kernel perspective challenges this notion. Since the intrinsic dimensionality of Imagenet is estimated to be $\sim 40$ [25], it is not unreasonable to expect that the span of desired features would be of similar dimensionality. It is, thus, intriguing that these frameworks mandate the use of an $\sim 8192-d$ projector head to capture the intricacies of the corresponding data augmentation kernel. This discrepancy can be explained by observing the learning dynamics of a linearized model under the BarlowTwins loss optimization [28]. These dynamics reveal that initializing the projection weight matrix in alignment with the eigenfunctions of the data kernel retains this alignment throughout the learning process. Notably, a high-dimensional projector is more likely to have a greater initialization span than its low-dimensional counterpart, increasing the likelihood of overlap with the relevant eigenfunctions. We hypothesize that it is possible to rectify this issue by using a stronger orthogonalization constraint for low-dimensional projectors, thereby rendering them sufficient for good feature learning.
4.3 Corollary 2: Multiple augmentations improve optimization
@@
Theorem 4.1 implies that the invariance loss optimization would ideally entail using the $T_{M}$ operator, thereby requiring many augmentations for each sample $x$. Using only two augmentations per sample yields a noisy estimate of $T_{M}$, yielding spurious eigenpairs [31] (see Appendix). These spurious eigenpairs add stochasticity to the learning dynamics and hinder the alignment of the learned features with the eigenfunctions of the data kernel [28]. We hypothesize that improving this estimation error by increasing the number of augmentations could alleviate this issue and improve the speed and quality of feature learning.
@@
Increasing the number of augmentations (say $m$) in BarlowTwins and VICReg comes with added compute costs. A straightforward approach would involve calculating the invariance loss for every pair of augmentations, resulting in $\mathcal{O}(m^{2})$ operations. However, Theorem 4.1 proposes an alternative method that uses the sample estimate of $T_{M}$, thereby requiring only $\mathcal{O}(m)$ operations. Both these strategies are functionally equivalent (see Appendix), but the latter is computationally more efficient. In summary, Theorem 4.1 establishes a mechanistic role for the number of data augmentations, paving the way for a computationally efficient multi-augmentation framework:
@@ 5 Experiments
In our experiments, we seek to serve two purposes: (i) provide empirical support for our theoretical insights and (ii) present practical primitives for designing efficient self-supervised learning routines. In summary, with extensive experiments across learning algorithms (BarlowTwins, VICReg) and training datasets (CIFAR-10/STL-10), we establish that
  • low-dimensional projectors as sufficient for learning good representations.
@@
  • multi-augmentation improves downstream accuracy and convergence.
@@
  • multi-Augmentation improves sample efficiency in SSL pretraining, i.e., recovering similar performance with significantly fewer unlabelled samples.
@@
Experiment Setup: We evaluate the effectiveness of different pretraining approaches for non-contrastive SSL algorithms using image classification as the downstream task. Across all experiments, we use linear probing with Resnet-50 as the feature encoder backbone. On CIFAR-10, all models are pretrained for 100 epochs, while STL-10 models are pretrained for 50 epochs. All runs are averaged over 3 seeds; error bars indicate standard deviation. Other details related to optimizers, learning rate, etc., are presented in the Appendix.
5.1 Sufficiency of Low-dimensional projectors
@@
Figure 3: Low-dimensional projectors are sufficient for good feature learning. We demonstrate that using a higher orthogonality constraint ($\lambda$ for D, F and $\lambda_{eff}=\frac{1}{d\lambda}$ for E) for lower projector dimensionality can achieve similar performance over a wide range of projector dimensions ($d$).
@@
Existing works recommend using high-dimensional MLPs as projectors (e.g., d=8192 for Imagenet in [32, 5]), and show significant degradation in performance for a fixed redundancy coefficient ($\lambda$). To reproduce this result, we run a grid search to find the optimal coefficient $(\lambda^{*}_{8192})$ for $d=8192$ and show that performance progressively degrades for lower $d$ if the same coefficient $\lambda^{*}_{8192}$ is reused for $d\in\{64,128,256,512,1024,2048,4096,8192\}$.
@@
Our insights in Section 4.2 suggest low-dimensional projectors should recover similar performance with appropriate orthogonalization. To test this, we find the best $\lambda$ by performing a grid search independently for each $d\in\{64,128,256,512,1024,2048,4096,8192\}$. Low-dimensional projectors are sufficient, as illustrated in Figure 3. Strikingly, we also observe that the optimal $\lambda_{d}\propto 1/d$ is in alignment with our theoretical insights.
@@
Recommendataions: Start with low-dimensional projector, using $\lambda=\mathcal{O}(\frac{1}{d})$, and sweep over $(pdim=d,\lambda=\mathcal{O}\left(\frac{1}{d}\right))$ if needed.
@@ 5.2 Multiple Augmentations Improve Performance and Convergence
Although some SSL pretraining approaches, like SWaV, incorporate more than two views, the most widely used heuristic in non-contrastive SSL algorithms involves using two views jointly encoded by a shared backbone. In line with this observation, our baselines for examining the role of multiple augmentations use two views for computing the cross-correlation matrix.
@@
Figure 4: Using multiple augmentations improves representation learning performance and convergence. (A-C) Across BarlowTwins and VICReg for CIFAR-10 and STL-10 pretraining, using 4 augmentations instead of 2 helps improve performance. (D-F) Although the 4-augmentations take longer for each epoch, its performance still trumps the 2-augmentation version of the algorithm at the same wall clock time.
@@
To understand the role of multiple augmentations in pretraining in light of the augmentation-kernel interpretation, we propose Equation 12, which generalizes Barlow-Twins and VICReg to the multi-augmentation setting. In particular, for #$\#augs\in\{2,4,8\}$, we pretrain Resnet-50 with the generalized NC-SSL loss for 100 epochs on CIFAR-10 and 50-epochs for STL-10. Building on the insight from the previous section, we use a 256-dimensional projector head for all experiments.
@@
In Figure 4, we track the downstream performance of the pre-trained models across training epochs, i.e., we extract features from intermediate checkpoints and train a linear classifier on top of the features. Here, we use the linear evaluation protocol as outlined by [6]. Figure 4(A-C) shows that pretraining with multiple augmentations outperforms the 2-augmentation baseline. Furthermore, we observe that the four-augmentation pre-trained models converge faster (both in terms of the number of epochs and wall-clock time) than their two-augmentation counterparts (see Figure 4(D-F)).
@@
Recommendatation: Using multiple augmentations ( $>2$) with the generalized NC-SSL loss is likely to improve convergence as well as downstream accuracy.
5.3 Sample Efficient Multi-View Learning
@@
Data Augmentation can be viewed as a form of data inflation, where the number of training samples is increased by $k$ (for $k$ augmentations). In this section, we examine the role of multi-augmentation in improving sample efficiency. In particular, we are interested in understanding if the same performance can be achieved with a fraction of the pretraining dataset.
@@
Figure 5: Multi-augmentation improves sample efficiency, recovering similar performance with significantly fewer unique samples in the pretraining dataset. Across BarlowTwins and VICReg pretraining on CIFAR-10 and STL-10, for the same effective dataset size (#$\#augs\times\#unique\_samples$), using more patches improves performance at the same epoch (A-C) or wall clock time (D-F). However, a tradeoff exists wherein doing more data augmentations fails to improve performance in the scarce data regime.
@@
To examine the relation between the number of augmentations and sample efficiency, we fixed the effective size of the inflated dataset. This is achieved by varying the fraction of the unique samples in the pretraining dataset depending on the number of augmentations $k\in\{2,4,8\}$, e.g., we use 1/2 the dataset for 4 views. We then evaluate the performance of the pre-trained models on the downstream task, where the linear classifier is trained on the same set of labeled samples. Strikingly, Figure 5 shows that using multiple augmentations can achieve similar (sometimes even better) performance with lesser pretraining samples, thereby indicating that more data augmentations can be used to compensate for smaller pretraining datasets.
@@ 6 Discussion
Summary: Our work presents a fresh theoretical analysis that sheds light on the implicit bias of non-contrastive SSL algorithms. We use these insights to unravel the impact of crucial design heuristics and offer practical recommendations that improve convergence while maintaining accuracy (on CIFAR-10/STL-10). We also show that the multi-augmentation framework can be used to learn good features from fewer unique samples in the pretraining dataset simply by improving the estimation of the data augmentation kernel.
@@
Open Questions: Looking ahead, it would be exciting to extend this analysis to other categories of SSL algorithms, such as Masked AutoEncoders (MAE). Furthermore, our insights provide opportunities to explore sample-efficient methods that rely on less data, which is particularly important in critical domains such as medical imaging, where data is often scarce and expensive.
Limitations Our algorithm relies on multiple views of the same image to improve the estimation of the data-augmentation kernel. Although this approach does add some extra computational overhead, it significantly speeds up the learning process. We can explore the possibility of making the current design more computationally efficient to improve it further.

### v2

@@ 3.3 Takeaway 1: Low-dimensional projectors can yield good representations
Given the proximity of the formulation of Equation 9 to that of BarlowTwins and VICReg losses, we will leverage existing heuristics that have been shown to work in practice. As such, BarlowTwins and VICReg frameworks call for high-dimensional projectors while using a weak orthogonalization regularization to facilitate good feature learning. We know, from Theorem 3.1, that the eventual goal of these frameworks is to learn the eigenfunctions of the underlying data similarity graph. For example, since the intrinsic dimensionality of Imagenet is estimated to be $\sim 40$ [32], it is not unreasonable to expect that the span of desired features would be of similar dimensionality. It is, thus, intriguing that the current practice would suggest using an $\sim 8192$-dim projector head to capture the intricacies of the corresponding augmentation-defined data similarity kernel. This discrepancy can be explained by analyzing the learning dynamics, as in Theorem 3.2. Notably, a high-dimensional projector is likelier to have a greater initialization span than its low-dimensional counterpart, thereby increasing the alignment between $V(F_{0})$ and relevant eigenfunctions of $T_{k}$. We hypothesize that a stronger orthogonalization constraint for low-dimensional projectors can rectify this issue, reducing the redundancy in the network’s output space and rendering it sufficient for good feature learning.
3.4 Takeaway 2: Multiple augmentations improve kernel approximation
@@
By comparing the invariance criterion formulation in the standard BarlowTwins and VICReg losses to Equation 7, it can be inferred that current practices use a sample estimate of $T_{k}$. Using only two augmentations per sample yields a noisy estimate of $T_{k}$, yielding spurious eigenpairs [41] (see Appendix C). These spurious eigenpairs add stochasticity in the learning dynamics, and coupled with Theorem 3.2, increase the redundancy in the learned feature space [11]. We hypothesize that improving this estimation error by increasing the number of augmentations could alleviate this issue and improve the speed and quality of feature learning.
@@
Of course, increasing the number of augmentations ($m$) in the standard BarlowTwins and VICReg loss improves the estimate of $T_{k}$ but comes with added compute costs – a straightforward approach would involve calculating the invariance loss for every pair of augmentations, resulting in $\mathcal{O}(m^{2})$ operations. However, Theorem 3.1 proposes an alternative method that uses the sample estimate of $T_{M}$, thereby requiring only $\mathcal{O}(m)$ operations, and hence is computationally more efficient while yielding functionally equivalent features (see Appendix B). In summary, Theorem 3.1 establishes a mechanistic role for the number of data augmentations, paving the way for a computationally efficient multi-augmentation framework:
@@ 4 Experiments
In our experiments, we seek to (i) provide empirical support for our theoretical insights and (ii) present practical primitives for designing efficient SSL routines. Since our proposed loss function is closest to the formulation of BarlowTwins/VICReg, we present empirical evidence comparing our proposal to these baselines. In summary, with extensive experiments across learning algorithms (BarlowTwins & VICReg) and training datasets (CIFAR-10, STL-10 & Imagenet-100), we establish the following:
  • low-dimensional projectors can yield good representations.
@@
  • multi-augmentation improves downstream accuracy, as well as convergence rate.
@@
  • multi-augmentation improves sample efficiency in SSL pretraining, i.e., recovering similar performance with significantly fewer unique unlabelled samples.
@@
Experiment Setup: We evaluate the effectiveness of different pretraining approaches using image classification as the downstream task. Across all experiments, we pretrain a Resnet feature encoder backbone for 100 epochs (see Section E.1 for longer pretraining results) and use linear probing on the learned representations [^1] Code: https://github.com/kumarkrishna/fastssl. All runs are averaged over 3 seeds; error bars indicate standard deviation. Other details related to optimizers, learning rate, etc., are presented in the Appendix D.
4.1 Low-dimensional projectors can yield good representations
@@
Figure 2: Low-dimensional projectors can yield good representations. We demonstrate that using a higher orthogonality constraint, $\beta$, for lower projector dimensionality can achieve similar performance over a wide range of projector dimensions ($d$).
pdim Barlow Twins VICReg
fixed $\beta$ optimal $\beta^{*}$ fixed $\beta$ optimal $\beta^{*}$
64 $73.6\pm 0.9$ $82.1\pm 0.2$ $68.9\pm 0.2$ $81.9\pm 0.1$
256 $75.9\pm 0.7$ 83.4 $\pm$ 0.4 $75.3\pm 0.2$ $81.9\pm 0.3$
1024 $81.3\pm 1.0$ $82.9\pm 0.3$ $79.2\pm 0.9$ 82.5 $\pm$ 0.9
8192 $82.2\pm 0.4$ $82.2\pm 0.4$ $80.4\pm 1.5$ $80.4\pm 1.5$
Table 1: Optimizing for orthogonality appropriately allows low-dimensional projectors to match the performance (on CIFAR-10) of much higher-dimensional projectors.
@@
Existing works recommend using high-dimensional MLPs as projectors (e.g., d=8192 for Imagenet in [43, 5]), and show significant degradation in performance when using lower-dimensional projectors for a fixed redundancy coefficient ($\beta$). To reproduce this result, we run a grid search to find the optimal coefficient $(\beta^{*}_{8192})$ for $d=8192$ and show that performance progressively degrades for lower $d$ if the same coefficient $\beta^{*}_{8192}$ is reused for $d\in\{64,128,256,512,1024,2048,4096,8192\}$.
@@
Our insights in Section 3.3 suggest low-dimensional projectors should recover similar performance with appropriate orthogonalization. To test this, we find the best $\beta$ by performing a grid search independently for each $d\in\{64,128,256,512,1024,2048,4096,8192\}$. As illustrated in Figure 2, using low-dimensional projectors yield features with similar downstream task performance, compared to the features obtained using high-dimensional projectors. Strikingly, we also observe that the optimal $\beta_{d}\propto 1/d$, which aligns with our theoretical insights.
{mdframed}
@@
Recommendation: Start with low-dimensional projector, using $\beta=\mathcal{O}(\frac{1}{d})$, and sweep over $(pdim=d,\beta=\mathcal{O}\left(\frac{1}{d}\right))$ if needed.
@@ 4.2 Multiple Augmentations Improve Performance and Convergence
Although some SSL pretraining approaches, like SWaV [10], incorporate more than two views, the most widely used heuristic in non-contrastive SSL algorithms involves using two views jointly encoded by a shared backbone. In line with this observation, our baselines for examining the role of multiple augmentations use two views for computing the cross-correlation matrix.
@@
Figure 3: Using multiple augmentations improves representation learning performance and convergence. (A-C) Across BarlowTwins for CIFAR-10, STL-10 and Imagenet-100 pretraining, using 4 augmentations instead of 2 helps improve performance. Please see Section E.3 for more results.
augs pdim BarlowTwins VICReg
Time (min) Time (min)
2 8192 $99.36\pm 0.01$ $94.36\pm 0.01$
2 256 $62.34\pm 0.06$ $51.73\pm 0.04$
4 256 43.09 $\pm$ 0.20 39.02 $\pm$ 0.05
Table 2: Using multiple augmentations yields faster convergence, with reduced time to reach baseline performance on CIFAR-10, i.e. performance of feature encoder pretrained with an 8192-dim projector and 2 augmentations.
@@
To demonstrate the role of multiple augmentations in pretraining, we adapt the invariance criterion of BarlowTwins/VICReg to be in line with Equation 10. In particular, for #$\#augs\in\{2,4,8\}$, we pretrain a Resnet-50 encoder with our proposed loss. Building on the insight from the previous section, we use a 256-dimensional projector head for all multi-augmentation experiments.
@@
In Figure 3, we track the downstream performance of the pretrained models across training epochs. For performance evaluation, we use the linear evaluation protocol as outlined by [13]. Figure 3(A-C) shows that pretraining with multiple augmentations outperforms the 2-augmentation baseline. Furthermore, we observe that the four-augmentation pretrained models converge faster (both in terms of the number of epochs and wall-clock time) than their two-augmentation counterparts (see Figure 3(D-F)). Additionally, we show in Section E.2 that our framework can also be applied to multi-augmentation settings like SWaV, where not all augmentations are of the same resolution.
{mdframed}
@@
Recommendatation: Using multiple augmentations ( $>2$) is likely to improve convergence as well as downstream accuracy.
4.3 Sample Efficient Multi-augmentation Learning
@@
Data Augmentation can be viewed as a form of data inflation, where the number of training samples is increased by $k$ (for $k$ augmentations). In this section, we examine the role of multi-augmentation in improving sample efficiency. In particular, we are interested in understanding if the same performance can be achieved with a fraction of the pretraining dataset, simply by using more augmentations.
@@
Figure 4: Multi-augmentation improves sample efficiency, recovering similar performance with significantly fewer unique samples in the pretraining dataset. Across BarlowTwins pretraining on CIFAR-10, STL-10 and Imagenet-100 for the same effective dataset size (#$\#augs\times\#unique\_samples$), using more patches improves performance at the same epoch (A-C). However, a tradeoff exists wherein more data augmentations fail to improve performance in the scarce data regime.
augs pdim Percentage BarlowTwins VICReg
of Dataset Time (min) Time (min)
2 8192 100 % 63.43 $\pm$ 0.02 66.05 $\pm$ 0.01
2 256 100 % 39.52 $\pm$ 0.04 40.64 $\pm$ 0.04
4 256 50 % 28.25 $\pm$ 0.01 32.39 $\pm$ 0.01
8 256 25 % 27.74 $\pm$ 0.01 34.76 $\pm$ 0.01
Table 3: Time required to pass 80% accuracy on CIFAR-10 when pretraining on fraction of the dataset, while using multiple augmentations. See Figure 5 for further discussion.
@@
To examine the relation between the number of augmentations and sample efficiency, we fixed the effective size of the inflated dataset. This is achieved by varying the fraction of the unique samples in the pretraining dataset depending on the number of augmentations $k\in\{2,4,8\}$, e.g., we use 50% of the dataset for 4 views. We then evaluate the performance of the pretrained models on the downstream task, where the linear classifier is trained on the same set of labeled samples. Strikingly, Figure 4 shows that using multiple augmentations can achieve similar (sometimes even better) performance with lesser pretraining samples, thereby indicating that more data augmentations can be used for feature learning to compensate for smaller pretraining datasets.
{mdframed}
@@
5 Related Work
Self-Supervised Pretraining requires significant compute resources and most practitioners rely on empirical heuristics (see SSL cookbook [3] for a summary). While recent advances in SSL theory explore learning dynamics in linear (or shallow) models [39, 40], with a focus on understanding dimensionality collapse [20, 24], the theoretical underpinnings of most of the heuristics considered essential for good feature learning, are missing.
Contrastive SSL has received more theoretical attention, owing to its connection with metric learning and noise contrastive estimation [29, 4, 25]. In particular, HaoChen et al. [22] provide a theoretical framework for the SimCLR loss from an augmentation graph perspective, which leads to practical recommendations. Subsequently, Garrido et al. [19] establish a duality between contrastive and non-contrastive learning objectives, further bridging the gap between theory and practice.
Non-contrastive SSL algorithms’ theoretical foundations have received more attention recently [9, 44]. Prior works [2, 19, 18] have demonstrated that with modified learning objectives, low-dimensional projectors yield representations with good downstream performance. Similarly, previous works have demonstrated notable performance boosts when using a multi-patch framework in contrastive [17] and non-contrastive SSL [10, 42]. However, the theoretical basis for the benefits and trade-offs of either low-dimensional projectors or multiple augmentations is largely unclear. It is worth noting that Schaeffer et al. [34] present an information-theoretic perspective of the recently proposed non-contrastive SSL loss that leverages multiple augmentations, namely MMCR [42], but the computational advantages of using multiple augmentations on the learning dynamics is an active area of research.
Deep Learning theory has made significant strides in understanding the optimization landscape and dynamics of supervised learning [1]. In concurrent works [44, 9], the interplay between the inductive bias of data augmentations, architectures, and generalization has been explored from a purely theoretical perspective, establishing an equivalence among different SSL losses [44]. Furthermore, Simon et al. [36] used a more straightforward formulation of the BarlowTwins loss and investigated the learning dynamics in linearized models for the case when the invariance and orthogonalization losses have equal penalties. Although such a setting rarely used in practice, their approach serves as an inspiration for our work in studying the learning dynamics of non-contrastive SSL losses.
@@ 6 Discussion
Summary: Our work builds on existing theoretical results that establish an equivalence among different SSL frameworks, and proposes a compute-efficient reformulation of the common SSL loss. Using this loss reformulation and a study of the optimization dynamics, we proposed practical recommendations to improve the sample and compute efficiency of SSL algorithms. Specifically, we recommended low-dimensional projectors with increased orthogonality constraints and multi-augmentation frameworks, and we verified the effectiveness of these recommendations empirically. It is worth noting that our multi-augmentation formulation improves the efficiency of learning without altering the desiderata of SSL, i.e. the network learns the same feature space using our proposed multi-augmentation framework as with the original SSL formulation in the limit of infinite pretraining budget. To demonstrate this equivalence between the original SSL loss and our proposed version, we show in Section E.1 that longer pretraining on the 2-augmentation loss leads to similar downstream performance as the multi-augmentation versions (4 and 8 augmentations).
We also showed that the multi-augmentation framework can be used to learn good features from fewer unique samples in the pretraining dataset simply by improving the estimation of the data augmentation kernel. This result has direct implications on improving the Pareto frontier of samples-vs-performance for SSL pretraining, wherein we can achieve better downstream performance when limited number of samples are available in the pretraining dataset.
@@
Connections to Downstream performance: While our core theoretical results are aimed at accelerating convergence of the SSL loss itself, our empirical results highlight an improved downstream task performance earlier during pretraining. While this discrepancy might seem counter-intuitive at first, it is worth noting that the SSL loss inherent influences downstream performance as it encourages clustering of semantically similar images in the representation space. Such clustering properties in the representation space facilitates easier classification through methods k-nearest neighbors or linear decoding for a large number of tasks that rely on the semantic content of images. Previous works [20, 2, 18, 37] have discussed in detail how certain geometric properties of the learned representation space are connected to the linear classification performance for arbitrary decision boundaries, in expectation. However, an in-depth analysis of downstream tasks that are more amenable to linear decoding from the learned SSL representation space requires framing metrics of alignment between the pretraining objective (SSL desiderate) and the downstream task labels, and is an active area of research.
Open Questions: Looking ahead, it would be exciting to extend this analysis to other categories of SSL algorithms, such as Masked AutoEncoders (MAE). Furthermore, our insights provide opportunities to explore sample-efficient methods that rely on less data, which is particularly important in critical domains such as medical imaging, where data is often relatively scarce and expensive. On a different note, it is intriguing that animals often spend extended periods of time exploring novel objects, likely to gain multiple views of the object [6, 28]. Given the theoretical underpinnings of the computational benefits of multi-augmentation SSL outlined in our work, it would be exciting to develop models of biological learning that leverage these insights and enable sample-efficient continual learning in similar environments.
Limitations: Our algorithm relies on multiple augmentations of the same image to improve the estimation of the data-augmentation kernel. Though this approach speeds up the learning process, it also adds some extra computational overhead, which means that the impact of faster learning on wall-clock time is less than might be hoped for. One way to mitigate the effects of this limitation would be to scale up to a multi-GPU setting, since the computations for each augmentation can be run on a separate GPU in parallel. This could help ensure that the improved speed of learning directly translates to a significantly reduced wall-clock time for training.
Impact Statement: The goal of our work is to advance the general field of visual representation learning. Although there are potential downstream societal consequences of our work, we feel there are no direct consequences that must be specifically highlighted here.